Ratio and Proportion
2. A new virus is released on the internet; the administrator of a department's Local Area Network ( LAN) is given five minutes by a manager to estimate the impact. The administrator samples 15 of the PCs connected to the LAN, and finds that 9 are infected; use proportion to estimate the number of infected PCs if there are a total of 202 PCs connected to the LAN.
This problem has me totally confused! Again, I am not good a word problems. I never know where to start, I have been trying to work out the problem, but I don't know where to begin. Please help me!

Answer by stanbon(75887)

A new virus is released on the internet; the administrator of a department's Local Area Network ( LAN) is given five minutes by a manager to estimate the impact. The administrator samples 15 of the PCs connected to the LAN, and finds that 9 are infected; use proportion to estimate the number of infected PCs if there are a total of 202 PCs connected to the LAN.
------------------
Use a proportion:
[9 infected/15 sampled] = [x infected/all the PCs]
(9/15) = (x/202)
x = 202(9/15)
x = 202(3/5)
x = 121.2
Round down to get x = 121 infected PC's out of 202
-------------------
Cheers,
Stan H.